Two boats leave a ramp traveling in opposite directions. The second boat is 10 miles per hour faster than the first. After 3 hours they are 150 miles apart. Find the speeds of the boats?
Answer by rfer(16322)

x+x+10=150/3
2x=40
x=20 mph
x+10=30 mph